Four years ago, a father was 2 times as old as his son, now their combined ages equal 107 years. thus the father's present age is what?
Answer by ankor@dixie-net.com(22740)

Four years ago, a father was 2 times as old as his son, now their combined ages equal 107 years.
thus the father's present age is what?
:
f = father's age
s = son's age
:
"Four years ago, a father was 2 times as old as his son,"
f - 4 = 2(s-4)
f - 4 = 2s - 8
f = 2s - 8 + 4
f = 2s - 4
:
" now their combined ages equal 107 years."
f + s = 107
f = (107-s)
replace f in the 1st equatin
107 - s = 2s - 4
107 + 4 = 2s + s
111 = 3s
s = 111/3
s = 37 yrs is son's age
then
107 - 37 = 70 is father's age
:
:
Check this in the statement:
"Four years ago, a father was 2 times as old as his son, "
70 - 4 = 2(37 - 4)
66 = 2(33)